OpenStudy (anonymous):

complete the square. -4x^2-16x+3=0

OpenStudy (anonymous):

i know that the answer is -4(x+2)^2+19 but i don't know how to get there..

OpenStudy (campbell_st):

oops rewrite it as \[-4x^2 -16x = -3\] is that ok..?

OpenStudy (anonymous):

yes and i know the next step would be \[-4(x ^{2}+4x)=-3\]

OpenStudy (campbell_st):

great... so what is needed inside the brackets to complete the square..?

OpenStudy (anonymous):

\[-4(x ^{2}+4x+4)=-3+4\]

OpenStudy (anonymous):

\[-4(x+2)^{2}=1\]

OpenStudy (campbell_st):

no you don't add 4.... you need to have \[-4(x^2 + 4x + 4) = -3 + (-4 \times 4)\] because of the common factor... think about what happens when you distribute

OpenStudy (anonymous):

.. i don't get that part

OpenStudy (campbell_st):

ok... 4 inside the brackets is correct... \[-4(x^3 + 6x + 4) = -3\] so I'll distribute \[-4x^2 -16x - 16 = -3 +??\] after distributing what value on the left needs to be added to the right...?

OpenStudy (anonymous):

OH.. 16.. because of the 4 that i factored out before.. so now i have to factor it back in

OpenStudy (campbell_st):

so the value has to be -16 so you get -4(x^2 + 4x + 4) = -3 -16 or -4(x + 2)^2 = -19 which becomes \[-4(x + 2)^2 + 19 = 0\]

OpenStudy (anonymous):

It feels great to finally understand how to complete the square after 3 years. thanks

OpenStudy (campbell_st):

lol.. well it is tough when the coefficient of x^2 isn't 1

OpenStudy (wolf1728):

This way might be easier: http://www.1728.org/quadr2.htm -4x^2 -16x +3=0 1) Move the "non X" term to the right: -4x^2 -16x= -3 2) Divide the equation by the coefficient of X² which in this case is -4 x^2 +4x = .75 3) Now here's the "completing the square" stage in which we: • take the coefficient of X which is +4 • divide it by 2 which equals 2 • square that number which equals 4 • then add it to both sides of the equation. x^2 +4x +4 = 4.75 (X+2) * (x+2) = 4.75 4) Take the square root of both sides of the equation: x +2 = square root (4.75) x +2 = 2.1794494718 AND x +2 = -2.1794494718 x = .1794494718 x = -4.1794494718